Consider a semiconducting material. What will be the energy of electrons and holes that constitute

excitons? Let m = rest mass of free electron, ε = dielectric constant = 15, Δ = 1.5 eV. Determine the

temperature at which excitons will be formed?


1
Expert's answer
2021-06-28T14:09:05-0400
"E(n)=-\frac{1}{n^2}\frac{\mu R_y}{m_0\epsilon_r^2}"

where "{\displaystyle {\text{Ry}}}" is the Rydberg unit of energy,

"{\displaystyle \varepsilon _{r}}"

is the relative permittivity,

"{\displaystyle \mu =\frac{(m_{e}^{*}m_{h}^{*})}{(m_{e}^{*}+m_{h}^{*})}}"

is the reduced mass of the electron and hole, and m0 is the electron mass.